# Interpolation in Classical Propositional Logic

### v1

@@ 2 Preliminaries
The set $\text{sub}(\varphi)$ of subformulas of a formula $\varphi$ is defined by induction by setting $\text{sub}(p)=\{p\}$, $\text{sub}(\top)=\{\top\}$, $\text{sub}(\bot)=\{\bot\}$, $\text{sub}(\varphi\wedge\psi)=\{\varphi\wedge\psi\}\cup\text{sub}(\varphi)\cup\text{sub}(\psi)$ and $\text{sub}(\neg\varphi)=\{\neg\varphi\}\cup\text{sub}(\varphi)$. Define the size $|\varphi|$ of a propositional formula $\varphi$ as the number of its distinct subformulas. This corresponds to representing formulas as directed acyclic graphs (dags) and is the standard measure used in circuit complexity theory. We also call $|\varphi|$ the dag-size of $\varphi$.
In contrast, if one takes into account the number of different occurrences of subformulas, one obtains the tree-size $s(\varphi)$ of $\varphi$, defined inductively by setting $s(p)=s(\top)=s(\bot)=1$, $s(\varphi\wedge\psi)=s(\varphi)+s(\psi)+1$, and $s(\neg\varphi)=s(\varphi)+1$. Clearly, $s(\varphi)\geq|\varphi|$ and it is conjectured that there is a superpolynomial (even exponential) gap between dag-size and tree-size in the sense that there are no functions $e$ mapping every propositional formula $\varphi$ to a logically equivalent propositional formula $e(\varphi)$ and polynomial function $f$ (that is, function of the form $f(n)=cn^{d}$ with $c,d$ natural numbers) such that $s(e(\varphi))\leq f(|\varphi|)$ for all $\varphi$. However, this conjecture remains open [42].
Central for interpolation is the notion of a signature: A signature $\sigma$ is a finite set of atoms. The signature $\text{sig}(\varphi)$ of a formula $\varphi$ is the set of atoms that occur in it. In other words, the signature of $\varphi$ are the non-logical symbols that occur in $\varphi$, whereas logical symbols such as $\wedge$, $\vee$, $\neg$, $\bot$ and $\top$ are not part of the signature of a formula. If $\mathit{sig}(\varphi)\subseteq\sigma$, we call $\varphi$ a $\sigma$-formula. Given two formulas $\varphi$ and $\psi$, we call $\psi$ a conservative extension of $\varphi$ if $\psi\models\varphi$, $\mathit{sig}(\varphi)\subseteq\mathit{sig}(\psi)$, and for every $\mathit{sig}(\varphi)$-formula $\chi$, $\psi\models\chi$ implies $\varphi\models\chi$.
@@ 3 Craig Interpolants
  • $\text{sig}(\chi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\psi)$.
@@
  • $\text{sig}(\chi)\subseteq\text{sig}(\varphi)\cap\text{sig}(\psi)$.
@@ 4 Uniform Interpolants
In this section we introduce a special form of Craig interpolants called uniform interpolants. While Craig interpolants are defined in terms of a pair of input formulas $\varphi,\psi$ with $\varphi\models\psi$, uniform interpolants are defined in terms of a single formula $\varphi$ and a signature $\sigma\subseteq\text{sig}(\varphi)$.
@@
Let $\varphi$ be a propositional formula and $\sigma\subseteq\text{sig}(\varphi)$. Then a propositional formula $\chi$ is called a uniform $\sigma$-interpolant for $\varphi$ if
@@
  • $\text{sig}(\chi)\subseteq\sigma$; and
  • for every propositional formula $\psi$, if $\varphi\models\psi$ and $\text{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\chi\models\psi$. [^2] Uniform interpolants are not always defined in exactly the same way. For instance, instead of starting from the signature $\sigma$ of the uniform interpolant, one can start from the signature $\text{sig}(\varphi)\setminus\sigma$ that is ‘forgotten’, or the condition $\text{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$ is replaced by the condition $\text{sig}(\psi)\subseteq\sigma$. In the context of propositional logic these different definitions are equivalent.
Observe that, given $\varphi$ and $\sigma\subseteq\text{sig}(\varphi)$, a uniform $\sigma$-interpolant for $\varphi$ is a Craig interpolant for all pairs $\varphi,\psi$ such that $\varphi\models\psi$ and $\psi$ shares with $\varphi$ only atoms that are in $\sigma$. In contrast to Craig interpolants, uniform interpolants are uniquely determined (up to logical equivalence) by $\varphi$ and $\sigma$. In what follows, we therefore refer to the uniform $\sigma$-interpolant for a formula $\varphi$.
@@
Note that in Example 4.11, the uniform $\sigma$-interpolant is always the logically strongest Craig interpolant for $\varphi,\psi$. This is no accident. To show this in general, assume that $\chi$ is the logically strongest Craig interpolant for $\varphi,\psi$. Let $\sigma=\text{sig}(\varphi)\cap\text{sig}(\psi)$ and let $\chi^{\prime}$ be the uniform $\sigma$-interpolant for $\varphi$. We show that $\chi$ and $\chi^{\prime}$ are logically equivalent. Clearly $\chi^{\prime}\models\chi$ since $\varphi\models\chi$ and $\text{sig}(\chi)\subseteq\sigma$. We also have $\chi\models\chi^{\prime}$ since $\chi^{\prime}$ is a Craig interpolant for $\varphi,\psi$ and $\chi$ is the logically strongest one.
@@
Let $\varphi$ be a propositional formula and $\sigma\subseteq\text{sig}(\varphi)$. Then a propositional formula $\chi$ is called a left-uniform $\sigma$-interpolant for $\varphi$ if
@@
  • $\text{sig}(\chi)\subseteq\sigma$; and
  • for every propositional formula $\psi$, if $\psi\models\varphi$ and $\text{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\psi\models\chi$.
@@
In propositional logic, if $\varphi$ is a propositional formula and $\sigma\subseteq\text{sig}(\varphi)$, then a uniform $\sigma$-interpolant for $\varphi$ exists.
@@ 5 Beth Definability
Let $\varphi$ be a propositional formula, $\sigma$ be a set of atoms, and $p\not\in\sigma$. Then a propositional formula $\psi$ is a $\sigma$-definition of $p$ under $\varphi$ if $\text{sig}(\psi)\subseteq\sigma$ and $\varphi\models p\leftrightarrow\psi$. $p$ is called explicitly $\sigma$-definable under $\varphi$ if there exists a $\sigma$-definition of $p$ under $\varphi$.
@@
Any Craig interpolant $\chi$ of $\varphi\wedge p,\varphi^{\prime}\rightarrow p^{\prime}$ is a $\sigma$-definition of $p$ under $\varphi$.
@@
The special case of the projective Beth definability property where $\sigma$ is not a parameter but defined as $\sigma=\text{sig}(\varphi)\setminus\{p\}$ is known as the Beth definability property. Note also that the notions of explicit and implicit definability, and the projective Beth definability property can be easily generalized to definitions of compound propositional formulas in place of atoms $p$.
@@
  • $\chi$ is an explicit $\text{sig}(\varphi)\cap\text{sig}(\psi)$-definition of $\psi$ under $\psi\rightarrow\varphi$.
@@ 6 Generalisations and an Application to Theory Decomposition
Craig interpolation has been generalised in many different ways. In particular, in applications of Craig interpolation to program verification various inductive generalisations have been introduced, including inductive sequences of interpolants [19, 36], tree interpolants [18, 7], and disjunctive interpolation [43, 44], in increasing order of generality, we refer the reader to [51] for further discussion. In all these generalisations, one considers multiple formulas (for instance, as labels of nodes in a tree), and the interpolants capture the interaction between these formulas in their shared signature.
@@
  • $\text{sig}(\chi_{i})\subseteq\text{sig}(\varphi_{i})\cap\bigcup_{j\in\{1,\ldots,l\},j\neq i}\text{sig}(\varphi_{j})$ for $1\leq i\leq l$
@@
In applications to program verification, these interpolants have been studied for $l=k+1$ and $\varphi_{k+1}=\bot$ [34, Sect. 5], where they are called symmetric interpolants. We focus on a brief discussion of another special case of Theorem 6.20 called parallel interpolation. In what follows, $T$ denotes a finite set of propositional formulas which, intuitively, axiomatises some theory or knowledge base of interest. We often identify $T$ with the conjunction $\bigwedge_{\chi\in T}\chi$; for example, $T\models\varphi$ stands for $\bigwedge_{\chi\in T}\chi\models\varphi$. Now, parallel interpolation is the special case in which $l=k+1$ and $\text{sig}(\varphi_{i})\cap\text{sig}(\varphi_{j})=\emptyset$ for $1\leq i<j\leq k$. So, given $T=\{\varphi_{1},\ldots,\varphi_{n}\}$ with $\text{sig}(\varphi_{i})\cap\text{sig}(\varphi_{j})=\emptyset$ for $1\leq i<j\leq n$ and a formula $\psi$ with $T\models\psi$, there are formulas $\chi_{1},\ldots,\chi_{n}$, called parallel interpolants, with
@@
  • $\text{sig}(\chi_{i})\subseteq\text{sig}(\varphi_{i})\cap\text{sig}(\psi)$ for $1\leq i\leq n$; and
@@
Intuitively, the parallel interpolants $\chi_{i}$ capture, in the shared atoms of $\varphi_{i}$ and $\psi$, the contribution of $\varphi_{i}$ to the derivation of $\psi$ from $T$. Although fairly straightforward to derive from Craig interpolation, parallel interpolation has some unexpected applications. Here we use it to show that finest decompositions of theories into axioms with mutually disjoint signatures are uniquely determined (up to logical equivalence). In detail, let $T$ be a set of formulas and let $\sigma_{1},\ldots,\sigma_{n}$ be a partition of $\text{sig}(T)$. Then $\sigma_{1},\ldots,\sigma_{n}$ is called a splitting of $T$ if there are formulas $\chi_{1},\ldots,\chi_{n}$ such that $\text{sig}(\chi_{i})\subseteq\sigma_{i}$ for $1\leq i\leq n$ and $\{\chi_{1},\ldots,\chi_{n}\}$ axiomatises $T$, that is to say, $\bigwedge_{1\leq i\leq n}\chi_{i}$ and $\bigwedge_{\chi\in T}\chi$ are logically equivalent. Splittings formalise the idea of decomposing a theory $T$ into axioms that speak about mutually disjoint subject matters. We say that a partition $\sigma_{1},\ldots,\sigma_{n}$ of $\text{sig}(T)$ is at least as fine as partition $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ $\subseteq\text{sig}(T)$ if for all $\sigma_{i}^{\prime}$ there exists $\sigma_{j}$ with $\sigma_{j}\subseteq\sigma_{i}^{\prime}$.
@@
Assume splittings $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ of $T$ are given. We show that the set of non-empty $\sigma_{i}\cap\sigma_{j}^{\prime}$, with $1\leq i\leq n$ and $1\leq j\leq m$ are again a splitting of $T$. The claim then follows directly by assuming that $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ are finest splittings of $T$. Let $\varphi_{1},\ldots,\varphi_{n}$ and $\psi_{1},\ldots,\psi_{m}$ be axiomatisations of $T$ witnessing that $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ are splittings of $T$, respectively. Then take the parallel interpolants $\chi_{ij}$ for the entailments $\{\varphi_{1},\ldots,\varphi_{n}\}\models\psi_{j}$ with $\varphi_{i}\models\chi_{ij}$, for $1\leq i\leq n$ and $1\leq j\leq m$. Then $\text{sig}(\chi_{ij})\subseteq\sigma_{i}\cap\sigma_{j}^{\prime}$ and the $\chi_{ij}$ axiomatise $T$. So they witness that the non-empty $\sigma_{i}\cap\sigma_{j}^{\prime}$ are a splitting of $T$.

### v2

@@ 2 Preliminaries
The set $\mathit{sub}(\varphi)$ of subformulas of a formula $\varphi$ is defined by induction by setting $\mathit{sub}(p)=\{p\}$, $\mathit{sub}(\top)=\{\top\}$, $\mathit{sub}(\bot)=\{\bot\}$, $\mathit{sub}(\varphi\wedge\psi)=\{\varphi\wedge\psi\}\cup\mathit{sub}(\varphi)\cup\mathit{sub}(\psi)$, $\mathit{sub}(\varphi\vee\psi)=\{\varphi\vee\psi\}\cup\mathit{sub}(\varphi)\cup\mathit{sub}(\psi)$, and $\mathit{sub}(\neg\varphi)=\{\neg\varphi\}\cup\mathit{sub}(\varphi)$. Define the size $|\varphi|$ of a propositional formula $\varphi$ as the number of its distinct subformulas. This corresponds to representing formulas as directed acyclic graphs (dags) and is the standard measure used in circuit complexity theory. We also call $|\varphi|$ the dag-size of $\varphi$.
In contrast, if one takes into account the number of different occurrences of subformulas, one obtains the tree-size $s(\varphi)$ of $\varphi$, defined inductively by setting $s(p)=s(\top)=s(\bot)=1$, $s(\varphi\wedge\psi)=s(\varphi)+s(\psi)+1$, $s(\varphi\vee\psi)=s(\varphi)+s(\psi)+1$, and $s(\neg\varphi)=s(\varphi)+1$. Clearly, $s(\varphi)\geq|\varphi|$ and it is conjectured that there is a superpolynomial (even exponential) gap between dag-size and tree-size in the sense that there are no functions $e$ mapping every propositional formula $\varphi$ to a logically equivalent propositional formula $e(\varphi)$ and polynomial function $f$ (that is, function of the form $f(n)=cn^{d}$ with $c,d$ natural numbers) such that $s(e(\varphi))\leq f(|\varphi|)$ for all $\varphi$. However, this conjecture remains open [42].
Central for interpolation is the notion of a signature: A signature $\sigma$ is a finite set of atoms. The signature $\mathit{sig}(\varphi)$ of a formula $\varphi$ is the set of atoms that occur in it. In other words, the signature of $\varphi$ are the non-logical symbols that occur in $\varphi$, whereas logical symbols such as $\wedge$, $\vee$, $\neg$, $\bot$ and $\top$ are not part of the signature of a formula. If $\mathit{sig}(\varphi)\subseteq\sigma$, we call $\varphi$ a $\sigma$-formula. Given two formulas $\varphi$ and $\psi$, we call $\psi$ a conservative extension of $\varphi$ if $\psi\models\varphi$, $\mathit{sig}(\varphi)\subseteq\mathit{sig}(\psi)$, and for every $\mathit{sig}(\varphi)$-formula $\chi$, $\psi\models\chi$ implies $\varphi\models\chi$.
@@ 3 Craig Interpolants
  • $\mathit{sig}(\chi)\subseteq\mathit{sig}(\varphi)\cap\mathit{sig}(\psi)$.
@@
  • $\mathit{sig}(\chi)\subseteq\mathit{sig}(\varphi)\cap\mathit{sig}(\psi)$.
@@ 4 Uniform Interpolants
In this section we introduce a special form of Craig interpolant called a uniform interpolant. While Craig interpolants are defined in terms of a pair of input formulas $\varphi,\psi$ with $\varphi\models\psi$, uniform interpolants are defined in terms of a single formula $\varphi$ and a signature $\sigma\subseteq\mathit{sig}(\varphi)$.
@@
Let $\varphi$ be a propositional formula and $\sigma\subseteq\mathit{sig}(\varphi)$. Then a propositional formula $\chi$ is called a uniform $\sigma$-interpolant for $\varphi$ if
@@
  • $\mathit{sig}(\chi)\subseteq\sigma$; and
  • for every propositional formula $\psi$, if $\varphi\models\psi$ and $\mathit{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\chi\models\psi$. [^2] Uniform interpolants are not always defined in exactly the same way. For instance, instead of starting from the signature $\sigma$ of the uniform interpolant, one can start from the signature $\mathit{sig}(\varphi)\setminus\sigma$ that is ‘forgotten’, or the condition $\mathit{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$ is replaced by the condition $\mathit{sig}(\psi)\subseteq\sigma$. In the context of propositional logic these different definitions are equivalent.
Observe that, given $\varphi$ and $\sigma\subseteq\mathit{sig}(\varphi)$, a uniform $\sigma$-interpolant for $\varphi$ is a Craig interpolant for all pairs $\varphi,\psi$ such that $\varphi\models\psi$ and $\sigma=\mathit{sig}(\psi)\cap\mathit{sig}(\varphi)$. In contrast to Craig interpolants, uniform interpolants are uniquely determined (up to logical equivalence) by $\varphi$ and $\sigma$. In what follows, we therefore refer to the uniform $\sigma$-interpolant for a formula $\varphi$.
@@
Note that in Example 4.11, the uniform $\sigma$-interpolant is always the logically strongest Craig interpolant for $\varphi,\psi$. This is no accident. To show this in general, assume that $\chi$ is the logically strongest Craig interpolant for $\varphi,\psi$. Let $\sigma=\mathit{sig}(\varphi)\cap\mathit{sig}(\psi)$ and let $\chi^{\prime}$ be the uniform $\sigma$-interpolant for $\varphi$. We show that $\chi$ and $\chi^{\prime}$ are logically equivalent. Clearly $\chi^{\prime}\models\chi$ since $\varphi\models\chi$ and $\mathit{sig}(\chi)\subseteq\sigma$. We also have $\chi\models\chi^{\prime}$ since $\chi^{\prime}$ is a Craig interpolant for $\varphi,\psi$ and $\chi$ is the logically strongest one.
@@
Let $\varphi$ be a propositional formula and $\sigma\subseteq\mathit{sig}(\varphi)$. Then a propositional formula $\chi$ is called a left-uniform $\sigma$-interpolant for $\varphi$ if
@@
  • $\mathit{sig}(\chi)\subseteq\sigma$; and
  • for every propositional formula $\psi$, if $\psi\models\varphi$ and $\mathit{sig}(\psi)\cap\mathit{sig}(\varphi)\subseteq\sigma$, then $\psi\models\chi$.
@@
In propositional logic, if $\varphi$ is a propositional formula and $\sigma\subseteq\mathit{sig}(\varphi)$, then a uniform $\sigma$-interpolant for $\varphi$ exists.
@@ 5 Beth Definability
Let $\varphi$ be a propositional formula, $\sigma$ be a set of atoms, and $p\not\in\sigma$. Then a propositional formula $\psi$ is a $\sigma$-definition of $p$ under $\varphi$ if $\mathit{sig}(\psi)\subseteq\sigma$ and $\varphi\models p\leftrightarrow\psi$. $p$ is called explicitly $\sigma$-definable under $\varphi$ if there exists a $\sigma$-definition of $p$ under $\varphi$.
@@
Let $\chi$ be any Craig interpolant for $\varphi\wedge p,\varphi^{\prime}\rightarrow p^{\prime}$. We show that $\chi$ is a $\sigma$-definition of $p$ under $\varphi$. By definition of Craig interpolants, $\mathit{sig}(\chi)\subseteq\sigma$. From $\varphi\wedge p\models\chi$ we obtain $\varphi\models p\rightarrow\chi$. Conversely, from $\chi\models\varphi^{\prime}\rightarrow p^{\prime}$ we obtain $\varphi^{\prime}\models\chi\rightarrow p^{\prime}$. Then, since $\mathit{sig}(\chi)\subseteq\sigma$, we obtain $\varphi\models\chi\rightarrow p$, as required.
@@
The special case of the projective Beth definability property where $\sigma$ is not a parameter but defined as $\sigma=\mathit{sig}(\varphi)\setminus\{p\}$ is known as the Beth definability property. Note also that the notions of explicit and implicit definability, and the projective Beth definability property can be easily generalized to definitions of compound propositional formulas in place of atoms $p$.
@@
  • $\chi$ is an explicit $\mathit{sig}(\varphi)\cap\mathit{sig}(\psi)$-definition of $\psi$ under $\psi\rightarrow\varphi$.
@@ 6 Generalisations and an Application to Theory Decomposition
Craig interpolation has been generalised in many different ways. In particular, in applications of Craig interpolation to program verification various inductive generalisations have been introduced, including inductive sequences of interpolants [20, 36], tree interpolants [19, 7], and disjunctive interpolation [43, 44], in increasing order of generality. We refer the reader to [51] for further discussion. In all these generalisations, one considers multiple formulas (for instance, as labels of nodes in a tree), and the interpolants capture the interaction between these formulas in their shared signature.
@@
  • $\mathit{sig}(\chi_{i})\subseteq\mathit{sig}(\varphi_{i})\cap\bigcup_{j\in\{1,\ldots,l\},j\neq i}\mathit{sig}(\varphi_{j})$ for $1\leq i\leq l$
@@
In applications to program verification, these interpolants have been studied for $l=k+1$ and $\varphi_{k+1}=\bot$ [34, Sect. 5], where they are called symmetric interpolants. We focus on a brief discussion of another special case of Theorem 6.20 called parallel interpolation. In what follows, $T$ denotes a finite set of propositional formulas which, intuitively, axiomatises some theory or knowledge base of interest. We often identify $T$ with the conjunction $\bigwedge_{\chi\in T}\chi$; for example, $T\models\varphi$ stands for $\bigwedge_{\chi\in T}\chi\models\varphi$. Now, parallel interpolation is the special case in which $l=k+1$ and $\mathit{sig}(\varphi_{i})\cap\mathit{sig}(\varphi_{j})=\emptyset$ for $1\leq i<j\leq k$. So, given $T=\{\varphi_{1},\ldots,\varphi_{n}\}$ with $\mathit{sig}(\varphi_{i})\cap\mathit{sig}(\varphi_{j})=\emptyset$ for $1\leq i<j\leq n$ and a formula $\psi$ with $T\models\psi$, there are formulas $\chi_{1},\ldots,\chi_{n}$, called parallel interpolants, with
@@
  • $\mathit{sig}(\chi_{i})\subseteq\mathit{sig}(\varphi_{i})\cap\mathit{sig}(\psi)$ for $1\leq i\leq n$; and
@@
Intuitively, the parallel interpolants $\chi_{i}$ capture, in the shared atoms of $\varphi_{i}$ and $\psi$, the contribution of $\varphi_{i}$ to the derivation of $\psi$ from $T$. Although fairly straightforward to derive from Craig interpolation, parallel interpolation has some unexpected applications. Here we use it to show that finest decompositions of theories into axioms with mutually disjoint signatures are uniquely determined (up to logical equivalence). In detail, let $T$ be a set of formulas and let $\sigma_{1},\ldots,\sigma_{n}$ be a partition of $\mathit{sig}(T)$. Then $\sigma_{1},\ldots,\sigma_{n}$ is called a splitting of $T$ if there are formulas $\chi_{1},\ldots,\chi_{n}$ such that $\mathit{sig}(\chi_{i})\subseteq\sigma_{i}$ for $1\leq i\leq n$ and $\{\chi_{1},\ldots,\chi_{n}\}$ axiomatises $T$, that is to say, $\bigwedge_{1\leq i\leq n}\chi_{i}$ and $\bigwedge_{\chi\in T}\chi$ are logically equivalent. Splittings formalise the idea of decomposing a theory $T$ into axioms that speak about mutually disjoint subject matters. We say that a partition $\sigma_{1},\ldots,\sigma_{n}$ of $\mathit{sig}(T)$ is at least as fine as partition $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ $\subseteq\mathit{sig}(T)$ if for all $\sigma_{i}^{\prime}$ there exists $\sigma_{j}$ with $\sigma_{j}\subseteq\sigma_{i}^{\prime}$.
@@
Assume splittings $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ of $T$ are given. We show that the set of non-empty $\sigma_{i}\cap\sigma_{j}^{\prime}$, with $1\leq i\leq n$ and $1\leq j\leq m$ are again a splitting of $T$. The claim then follows directly by assuming that $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ are finest splittings of $T$. Let $\varphi_{1},\ldots,\varphi_{n}$ and $\psi_{1},\ldots,\psi_{m}$ be axiomatisations of $T$ witnessing that $\sigma_{1},\ldots,\sigma_{n}$ and $\sigma_{1}^{\prime},\ldots,\sigma_{m}^{\prime}$ are splittings of $T$, respectively. Then take the parallel interpolants $\chi_{ij}$ for the entailments $\{\varphi_{1},\ldots,\varphi_{n}\}\models\psi_{j}$ with $\varphi_{i}\models\chi_{ij}$, for $1\leq i\leq n$ and $1\leq j\leq m$. Then $\mathit{sig}(\chi_{ij})\subseteq\sigma_{i}\cap\sigma_{j}^{\prime}$ and the $\chi_{ij}$ axiomatise $T$. So they witness that the non-empty $\sigma_{i}\cap\sigma_{j}^{\prime}$ are a splitting of $T$.